A soap manufacturer decides to spend 600000 rupees on radio, magazine and TV advertising. If he spends as
much on TV advertising as on magazines and radio together, and the amount spent on magazines and TV
combined equals five times that spent on radio. What is the amount to be spent on each type of advertising?

Answer by ikleyn(52788)

.

Let T = spending on TV;  R = spending on radio and M = spending on magazines.


From the condition, we have


    T + (R + M) = 600000

    T - (R + M) = 0.


It immediately implies that  


    T = R + M = 600000/2 = 300000.     (1)


In addition, from the condition, we are given that 

    T + M = 5R,    or

    300000 + M = 5R,    or

    5R - M = 300000.     (2)


So, from (1) and (2) we have


    R  + M = 300000,      (3)

    5R - M = 300000.      (4)


Add equations (3) and (4).  You will get

    6R     = 600000,

     R     = 600000/6 = 100000.


ANSWER.  R = 100000  (radio);  T = 300000 (TV)  and  M = 200000 (magazines).

Solved.